Find the value of k so that the point P lies on line L.
P(2,1), L: 3x + ky= 8

Answer by THANApHD(104)


3x+ky = 8
ky = 8 - 3x
y = 8/k - 3x/k ( standerd form of linear equation)

substitude the values
1= (8 - 3*2)/k
1= 2/K
K*1= 2
K = 2